ROTATIONAL SPECTRUM OF THE CO_{2}–CO_{2}–N_{2}O TRIMER

\underline{REBECCA A. APPLEMAN}, ROBERT L. KUCZKOWSKI, Department of Chemistry, University of Michigan, Ann Arbor, MI 48109-1055, USA.

The rotational spectrum of a trimer containing CO_{2} and N_{2}O has been assigned using a pulsed supersonic nozzle Fourier transform microwave spectrometer. 43 transitions of the normal species have been fit to a Watson A-reduction Hamiltonian to give the following constants: A = 1597.4630(29)MHz, B = 1232.9673(15)MHz, C = 831.1002(11)MHz, Δ_{J} = 2.610(33)kHz, Δ_{JK} = −3.63(14)kHz, Δ_{K} = 7.30(36)kHz, δ_{J} = 0.830(15)kHz, and δ_{K} = 0.71(10)kHz. These constants are consistent with predicted spectra of both CO_{2}–CO_{2}–N_{2}O and CO_{2}–N_{2}O–N_{2}O, but dipole moment data and preliminary isotopic work indicate that the observed spectrum probably belongs to the first species. The structure is believed to be barrel-like, with the three monomer units roughly parallel to each other. Theoretical modeling of the complex, including atom-atom electrostatic, dispersion and repulsion terms, gives minimum energy forms consistent with the observed rotational constants.